# Simulation of prosthetic vision with PRIMA system and enhancement of face representation

### v2

@@ 4 Discussion
In this paper, we presented an algorithm that simulates the loss of resolution and contrast in prosthetic vision as described by PRIMA patients. Results of this study illustrate why PRIMA patients can resolve letters but cannot perceive faces: unlike black-and-white letters, simulated prosthetic vision does not reproduce fine facial features, such as the eyebrows and lips.
Machine learning facial landmarking model helps identify the location and outline of the thinner facial features and make them programmatically thicker for any given image in real time. This procedure ensures that such features remain visible even in a low-resolution and low-contrast image of prosthetic vision. Similarly, understanding the specific characteristics of contrast sensitivity loss experienced by patients enables the application of the inverse of this function to the image before it is projected into the eye. While the accentuation of facial features can be done in the same manner for all patients, contrast enhancement will require a personalized clinical investigation of a patient’s contrast sensitivity loss function.
Integrating these or similar software enhancements into the PRIMA system should enable current patients to obtain their most wanted “upgrade” of face perception without any hardware or surgical changes. In addition, even though the next-generation implants with smaller and denser pixels should allow the representation of much higher spatial frequencies, the new devices will still require additional image processing for contrast enhancement in order to optimize face perception.
Acknowledgments

### v3

@@ 4 Discussion
In this paper, we presented an algorithm that simulates the loss of color, resolution and contrast in prosthetic vision, as described by PRIMA patients - ProViSim. Processing images with this simulator illustrates why PRIMA patients can resolve letters but cannot perceive faces: unlike black-and-white letters, simulated prosthetic vision does not reproduce fine facial features, such as the eyebrows and lips, or the complex greyscale of the various facial planes.
ML-based facial landmarking helps identify the location and outline of the thinner facial features and make them programmatically more pronounced for any given image in real time. This procedure ensures that such features remain visible even in a low-resolution and low-contrast image of prosthetic vision. Similarly, understanding the specific characteristics of contrast sensitivity loss experienced by patients will enable the application of the inverse of this function to the image before it is projected into the eye, thereby better retaining information on the facial structure.
Applying these enhancements on sighted individuals has significantly reduced their response time and improved their ability to recognize facial expressions and the underlying emotions. Moreover, if the function that governs contrast sensitivity loss is known, the color of the landmarking can be chosen such that it creates the thinnest and most natural overall output when perceived with the limitations imposed by prosthetic vision. This will require a personalized clinical investigation of a patient’s contrast sensitivity loss function and adjustment of their preferred enhancement settings.
For reading tasks, PRIMA patients have expressed a preference of viewing white letters on a black background, which effectively minimizes the percentage of the illuminated area in their prosthetic field of view. However, the same preference is unlikely to work for faces because they consist of several different planes and we are used to associate darkness with depth.
The findings of this study and the remaining open questions discussed here will be used as a foundation for a series of follow-up tests on PRIMA patients intended to measure their individual contrast sensitivity curves and examine their face perception with and without the suggested enhancements. Integrating these or similar software enhancements into the PRIMA system should enable current patients to obtain their most wanted “upgrade” of face perception without any hardware changes or surgical intervention. In addition, even though the next-generation implants with smaller pixels should allow the representation of much higher spatial frequencies, they will still require additional image processing for contrast enhancement in order to optimize the face perception.
Declarations
Ethics approval and consent to participate
This study was approved by the Stanford IRB panel on human subject research. Sixteen subjects recruited from Stanford University provided consent and participated in the study.
Consent for publication
Not applicable.
Availability of data and materials
Competing interests
Daniel Palanker: C (consultant) and P (patents licensed by Stanford University to Science Corp). All other authors have no conflicting interests to declare.
Funding
@@
Authors’ contributions
AKG, JP, and DP conceived the study. AKG, JP, YZ, and DP developed the methodology. JP and AKG implemented the computational framework. AKG collected the data. AKG and JP wrote the manuscript. JP, AKG and NJ prepared the figures. All authors reviewed and edited the manuscript.